OpenStudy (rachvanschaick):

A machinist turns the power on to a grinding wheel, which is at rest at time t = 0.00 s. The wheel accelerates uniformly for 10 s and reaches the operating angular velocity of 25 rad/s. The wheel is run at that angular velocity for 37 s and then power is shut off. The wheel decelerates uniformly at 1.5 rad/s2 until the wheel stops. In this situation, the time interval of angular deceleration (slowing down) is closest to: A) 17 s B) 15 s C) 19 s D) 21 s E) 23 s

OpenStudy (raffle_snaffle):

Use kinematics with angular velocity.

OpenStudy (shamim):

use angular acceleration =(final angular velocity -initial angular velocity) /time

OpenStudy (shamim):

anyway welcome to openstudy

OpenStudy (rachvanschaick):

But how?? Is it possible to see the problem worked out?

OpenStudy (irishboy123):

you will know the equations of motion for **linear** motion, eg one simple version is \(v = u + at\) \(x = u t + \frac{1}{2} a t^2\) \(v^2 = u^2 + 2 a x\) And if this were about a machine moving in a line I am sure you would have no problems For **rotary** motion, these equations have analogues: \(\omega_2 = \omega_1 + \alpha t\) \(\theta = \omega_1 t + \frac{1}{2} \alpha t^2\) \(\omega_2^2 = \omega_1^2 + 2 \alpha \theta\) look at the pattern matching, eg \(a \to \alpha\), to see what has happened. We're just switching letters around :-) the big idea is still motion with constant [here, angular] acceleration.